OpenStudy (anonymous):

Find the work done by the force F(x, y) = 2xyi + 4y2j acting along the piecewise smooth curve consisting of the line segments from (-2,2) to (0,0) and from (0,0) to (2,3).

OpenStudy (amistre64):

it looks to me kinda like you want the magnitude of force in the direction of the line segments given the results of the Force vector

OpenStudy (amistre64):



OpenStudy (amistre64):

looks to me like the magnitude along the line is equal to: F cos(a)

OpenStudy (amistre64):

\[F = \sqrt{(2xy)^2+(4y^2)^2}\] \[cos(\alpha) = \frac{a\cdot b}{|a|~|b|}\] \[cos(\alpha) = \frac{<2xy,4y^2>\cdot <1,-1>}{\sqrt{(2xy)^2+(4y^2)^2}~\sqrt2}\] \[F~cos(\alpha) = \frac{2xy-4y^2}{~\sqrt2}\] \[F~cos(\alpha) = \sqrt2~(xy-2y^2)\] \[Work=\int_{-2}^{0}\int_{2}^{0}F~cos(\alpha)~dy~dx\] maybe

OpenStudy (amistre64):

if its right, then the same concept for the other line segment